OpenStudy (anonymous):

(a^10m^24)^1/2(a^2m)^3 simplify

OpenStudy (anonymous):



OpenStudy (zzr0ck3r):

\((a^b)^c=a^{bc}\\(ab)^c=a^cb^c\)

OpenStudy (zzr0ck3r):

thats all you need to solve this

OpenStudy (anonymous):